### v1

@@ 4.4. Participants’ Approach and Task Limitations
Some exemplary images of finished models that participants created during the modeling sessions are depicted in Figure 7.
@@ 5. Discussion and Implications
In this section, we revisit the results of our quantitative and qualitative analysis in order to interpret them and derive implications for the future design of co-creative VR systems. Additionally, we discuss the generalizability and limitations of this work.
@@ 5.1. Highlighting does not alleviate the confusion caused by AI actions
We found the highlighting to decrease perceived partnership, enjoyment, and satisfaction with the finished model. At the same time, highlighting did not show effects on the predictability of the system, the perceived communication between humans and AI, the users’ attention toward the AI, or the clarity of AI intent, which we expected to be most likely affected by highlighting. On the qualitative side, the observational analysis also did not indicate that highlighting of object parts caused any differences in participant behavior. This is supported by participants never referring to highlighting during the modeling sessions or interviews. Taken together, our quantitative and qualitative findings suggest that highlighting is not as much of a mitigator against user confusion about AI intentions as expected and might even lead to a more negative overall experience. Future research should investigate whether these negative effects can be replicated, or if they occur similarly with different implementations of a highlighting variable.
lightgray Design implication for highlighting For co-creation use cases where users create one complete object at a time like in our use case, we recommend not using highlighting of changes. In particular, in combination with embodied AI avatars, it introduced more negative effects than it contributed to user satisfaction. Future work has to show if highlighting proves more useful for co-creation use cases of dynamic objects.
@@
Since the vast majority of participants watched all incremental visualization processes from beginning to end, it is not surprising that this variable significantly affected the attention participants paid to the AI. However, we also expected that the visualization aspect would affect the perceived efficiency of the system, which our quantitative analysis did not confirm. Considering that participants were likely aware that the AI was not generating model parts ”on the spot”, they likely interpreted this process as the loading time of the overall setup. This interpretation also explains why the incremental visualization did not show effects on predictability, competence, unusualness, value, creativity, or contribution, as we hypothesized before.
@@
Further, this lack of impact on the perception of the AI and its outputs could be caused by the specific incremental visualization we used. This incremental visualization was inspired by AI drawing agents that build images stroke-by-stroke. Transferring this concept directly into 3D space resulted in an AI agent that slowly draws individual horizontal lines of different colored pixels, which does not correspond to our experiences in the physical world. A possibly more appropriate 3D equivalent could be the molding or carving of primitives. Future work should investigate whether alternative ways of incrementally visualizing 3D modifications indeed have different effects on observers’ perceptions of the system.
@@
Another surprising finding is that incremental visualization increased the perceived unexpectedness of AI outputs. This contrasts with our hypothesis that incremental visualization would lead to a decrease in unexpectedness rather than an increase, as the longer visualization duration gave participants more time to familiarize themselves with the modified version of an object part. We attribute this reverse effect to participants having had more time to engage with the version of an object part due to the extra time (e.g., to think about what the fully instantiated version of the object part would look like or what kind of change they would have expected instead at this point). As a result, they perceived the AI outputs as more unexpected than in conditions where the output appeared immediately, and participants have no choice but to immediately accept them as the new version of an object part.
@@
Additionally, our interviews revealed that incremental visualization tended to be perceived more positively when an avatar appeared to carry out the incremental process. We attribute this to a more sympathetic attitude towards an AI with similar characteristics to the user. An AI that does not have a body, be it physical or virtual, might be expected not to be limited by time and space to the same degree as a human and might, therefore, seem more capable of performing actions that would be impossible for a human user. An embodied AI, on the other hand, might appear to be constrained by its surroundings in similar ways to a human. It would make sense for a human user to be more sympathetic toward AI that ”has to” slowly work on its task until it is finished, and therefore to feel more like the AI is equal to the user. One participant also applied this idea to the ”tool” that the AI used in the incremental process, which was a line that looked similar to the pointer ray that participants used to interact with the VR environment. It would be interesting for future work to study whether the effects of this similarity of tools only apply when a similarity of processes (i.e., an incremental process) is also given or whether the same effect can be observed given an instantaneous visualization.
@@
lightgray Design implication for incremental visualization We recommend using an immedeate creation of content for the co-creation of complete objects, where the process of creation is not important. We ground this in our findings where incremental generation did not reduce participants’ feelings of unexpectedness, while at the same time, the visualization took longer. However, we recommend incremental visualization for use cases where the process of creating objects is important and where users should observe the generation more closely and time is no relevant factor.
@@ 5.3. Embodiment not only affects the relationship to the AI but also to the 3D model created
Embodiment was the visualization aspect that participants had the strongest feelings about. The divergence in comfort levels in regard to the presence of the avatar was unexpected: To avoid an avatar design that created feelings of uncanniness in users (Mori et al., 2012), we intentionally chose an avatar for our AI that was heavily caricatured and only distantly resembled the form of a human, as well as one that had a friendly, child-like appearance. We do not know whether participants’ discomfort with the avatar was indeed a case of the uncanny valley, whether it was only due to the surprise participants felt when they were faced with the avatar for the first time, or whether it can be attributed to characteristics that are specific to co-creative systems.
@@
The divergence in participants’ discomfort with the avatar is also a possible explanation for why embodiment did not significantly affect ratings of factors measuring appeal. Nevertheless, our qualitative analysis indicated an improvement in the perceived relationship between users and the AI when the avatar was present, manifested by an increased tangibility of the relationship, the feeling of being less alone, and more empathy expressed towards the AI. This is consistent with quantitative data pointing to an effect of embodiment on partnership and AI-human communication.
@@
As hypothesized, AI embodiment did affect the extent to which participants felt that the AI contributed to the finished 3D model. This suggests that users’ perceptions of who created an artifact can be manipulated by factors external to the AI functionality itself. This finding is noteworthy considering currently ongoing discussions around ownership of AI-generated content, as the question of who created an artifact is often one of the most influential ones in deciding who owns it (Rezwana and Maher, 2023). Rezwana and Maher (Rezwana and Maher, 2023) found that this decision of ownership, and additionally users’ stances on other AI-related moral dilemmas, is also greatly influenced by the perception of a co-creative AI partner as either a collaborator or as a tool. Considering our finding that the embodiment of our AI significantly affected participants’ perceptions of the AI as a partner rather than a tool, we argue that such ethical questions should be kept in mind when deciding whether to integrate an AI avatar into a co-creative system. This is in line with Buschek et al. (Buschek et al., 2021) listing unclarity of ownership as a potential pitfall in the design of co-creative systems.
@@
lightgray Design implication for embodiment We recommend using an embodied AI for most co-creation use cases, as it increased the perceived support of the system, led to a higher perceived partnership and communication, and led to users paying more attention to the creation. However, users also felt reduced ownership of the final outcome, which should be considered when designing a system. For use cases where users just need a tool and a creative collaborative aspect is less relevant, the non-embodied AI might be the appropriate choice, as it is faster and users have a higher ownership of the final outcome.
@@ 6. Limitations and Future Work
Our work provides valuable insights and implications for the future design of co-creative systems for object generation in VR. In particular, we are confident that our technology agnostic approach, which does not rely on current AI capabilities or technologies, will ensure that the results remain relevant for future systems, even if these are not constrained by the same technological limitations.
However, we acknowledge that our experiment focused on specific AI representation modes, and other variants or visualizations may yield different results. We selected these specific modes of representation because of their importance for human collaboration and to explore their impact on users’ perceptions of the co-creative process and to stimulate further research and alternative implementations.
@@
Also, the limited number of possible AI modifications, which remained constant during iterative modifications, could have led participants to the conclusion that their AI partner did not actually generate object parts, but retrieved them from previously modeled parts. We hypothesize that this may have had an impact on the perception of the AI and its results as nothing about the AI’s capabilities and its ”generative” process changed between conditions.
@@
Since we performend the study as a controlled experiment, we acknowledge limitations in external validity in favor of establishing cause-effect relationships to recognize the impacts of individual representation modes.
There are several interesting directions for future work: One important area is exploring how different methods of emphasizing the AI’s contribution process might positively impact a user’s experience with a co-creative system, including whether such methods can enhance user navigation and coordination during collaboration. Additionally, investigating various ways to incrementally visualize the AI’s generation of a 3D object could reveal how these visualizations affect users’ perceptions of the AI and its creative process. There is also a need to examine the specific characteristics of avatars for co-creative AI that may cause discomfort to users, as well as how the perceived closeness to AI agents is affected by their embodiment, and which other factors might influence this perception. Finally, future research should also explore whether manipulating the visual appearance of an AI’s embodiment can alter people’s understanding and evaluation of the AI and its creations.
@@ 7. Conclusion
Research on co-creative systems has predominantly focused on how humans participate in the co-creative process, often overlooking how AI contributions are represented and perceived by users. This paper addresses this gap by examining how an AI’s contributions can be effectively represented for co-creative object generation in VR. Through a Wizard-of-Oz study in a VR-based co-creative 3D object-generating environment, we investigated the effects on the co-creative collaboration experience for three key representation modes that AI can use to communicate to the user: highlighting, incremental visualization, and embodiment of AI contributions. Finally, we derive implications for the design of tools for co-creative object-building: First, designers should avoid combining highlighting with embodiment to increase user satisfaction. Second, while an embodiment increases perceived support, it also reduces ownership of the created objects. Third, designers should favor immediate visualization when the creation process is secondary.

### v3

@@ 4.4. Participants’ Approach and Task Limitations
Some exemplary images of finished models created by participants during the modeling sessions are depicted in Figure 7.
@@ 5. Discussion and Implications
In this section, we revisit our quantitative and qualitative analysis results to interpret them and derive implications for the future design of co-creative VR systems. Additionally, we discuss the generalizability and limitations of this work.
@@ 5.1. Highlighting does not alleviate the confusion caused by AI actions
We found the highlighting to decrease perceived partnership, enjoyment, and satisfaction with the finished model. At the same time, highlighting did not affect the system’s predictability, the perceived communication between humans and AI, the users’ attention toward the AI, or the clarity of AI intent, which we expected to be most likely affected by highlighting. On the qualitative side, the observational analysis also did not indicate that highlighting object parts caused differences in participant behavior. This result is supported by participants never referring to highlighting during the modeling sessions or interviews. Our quantitative and qualitative findings suggest that highlighting is not as much of a mitigator against user confusion about AI intentions as expected and might lead to a more negative overall experience. Future research should investigate whether these adverse effects can be replicated or if they occur similarly with different implementations of a highlighting variable.
lightgray Design implication for highlighting For co-creation use cases where users create one complete object at a time, like in our use case, we recommend not highlighting changes. In particular, in combination with embodied AI avatars, it introduced more negative effects than it contributed to user satisfaction. Future work has to show if highlighting proves more useful for co-creation use cases of dynamic objects.
@@
Since most participants watched all incremental visualization processes from beginning to end, it is unsurprising that this variable significantly affected the attention participants paid to the AI. However, we also expected that the visualization aspect would affect the perceived efficiency of the system, which our quantitative analysis did not confirm. Considering that participants were likely aware that the AI was not generating model parts ”on the spot,” they likely interpreted this process as the loading time of the overall setup. This interpretation also explains why the incremental visualization did not affect predictability, competence, unusualness, value, creativity, or contribution, as we hypothesized.
@@
Further, this lack of impact on the perception of the AI and its outputs could be caused by the specific incremental visualization we used. This incremental visualization was inspired by AI drawing agents that build images stroke-by-stroke. Transferring this concept directly into 3D space resulted in an AI agent that slowly draws individual horizontal lines of different colored pixels, which does not correspond to our experiences in the physical world. A possibly more appropriate 3D equivalent could be the molding or carving of primitives. Future work should investigate whether alternative ways of incrementally visualizing 3D modifications affect observers’ perceptions of the system differently.
@@
Another surprising finding is that incremental visualization increased the perceived unexpectedness of AI outputs. This finding contrasts with our hypothesis that incremental visualization would lead to a decrease in unexpectedness rather than an increase, as the longer visualization duration gave participants more time to familiarize themselves with the modified version of an object part. We attribute this reverse effect to participants having had more time to engage with the version of an object part due to the extra time (e.g., to think about what the fully instantiated version of the object part would look like or what kind of change they would have expected instead at this point). As a result, they perceived the AI outputs as more unexpected than in conditions where the output appeared immediately, and participants had no choice but to immediately accept them as the new version of an object part.
@@
Additionally, our interviews revealed that incremental visualization was perceived more positively when an avatar appeared to carry out the incremental process. We attribute this to a more sympathetic attitude towards an AI with characteristics similar to the user’s. An AI that does not have a physical or virtual body might be expected not to be limited by time and space to the same degree as a human and might, therefore, seem more capable of performing actions that would be impossible for a human user. On the other hand, an embodied AI might appear to be constrained by its surroundings in similar ways to a human. It would make sense for a human user to be more sympathetic toward AI, which ”has to” slowly work on its task until it is finished, and therefore, to feel more like the AI is equal to the user. One participant also applied this idea to the ”tool” that the AI used in the incremental process, which was a line that looked similar to the pointer ray that participants used to interact with the VR environment. It would be interesting for future work to study whether the effects of this similarity of tools only apply when a similarity of processes (i.e., an incremental process) is also given or whether the same effect can be observed given an instantaneous visualization.
@@
lightgray Design implication for incremental visualization We recommend using an immediate creation of content to co-create complete objects, where the creation process is unimportant. We ground this in our findings where incremental generation did not reduce participants’ feelings of unexpectedness, while at the same time, the visualization took longer. However, we recommend incremental visualization for use cases where the process of creating objects is important and where users should observe the generation more closely, and time is no relevant factor.
@@ 5.3. Embodiment not only affects the relationship to the AI but also to the 3D model created
Embodiment was the visualization aspect that participants had the most substantial feelings about. The divergence in comfort levels regarding the presence of the avatar was unexpected: To avoid an avatar design that created feelings of uncanniness in users (Mori et al., 2012), we intentionally chose an avatar for our AI that was heavily caricatured and only distantly resembled the form of a human, as well as one that had a friendly, child-like appearance. We do not know whether participants’ discomfort with the avatar was indeed a case of the uncanny valley, whether it was only due to the surprise participants felt when they were faced with the avatar for the first time, or whether it can be attributed to characteristics that are specific to co-creative systems.
@@
The divergence in participants’ discomfort with the avatar is also a possible explanation for why embodiment did not significantly affect ratings of factors measuring appeal. Nevertheless, our qualitative analysis indicated an improvement in the perceived relationship between users and the AI when the avatar was present, manifested by an increased tangibility of the relationship, the feeling of being less alone, and more empathy expressed towards the AI. This result is consistent with quantitative data pointing to an effect of embodiment on partnership and AI-human communication.
@@
As hypothesized, AI embodiment did affect the extent to which participants felt that the AI contributed to the finished 3D model. This result suggests that factors external to the AI functionality can manipulate users’ perceptions of who created an artifact. This finding is noteworthy considering ongoing discussions around ownership of AI-generated content, as the question of who created an artifact is often one of the most influential in deciding who owns it (Rezwana and Maher, 2023). Rezwana and Maher (Rezwana and Maher, 2023) found that this decision of ownership, and additionally users’ stances on other AI-related moral dilemmas, is also greatly influenced by the perception of a co-creative AI partner as either a collaborator or as a tool. Considering our finding that the embodiment of our AI significantly affected participants’ perceptions of the AI as a partner rather than a tool, we argue that such ethical questions should be kept in mind when deciding whether to integrate an AI avatar into a co-creative system. This result is in line with Buschek et al. (Buschek et al., 2021) listing unclarity of ownership as a potential pitfall in the design of co-creative systems.
@@
lightgray Design implication for embodiment We recommend using an embodied AI for most co-creation use cases, as it increases the perceived support of the system, leads to a higher perceived partnership and communication, and leads to users paying more attention to the creation. However, users also felt reduced ownership of the outcome, which should be considered when designing a system. For use cases where users need a tool and the creative, collaborative aspect is less relevant, the non-embodied AI might be the appropriate choice, as it is faster, and users have higher ownership of the outcome.
@@ 6. Limitations and Future Work
Our work provides valuable insights and implications for the future design of co-creative systems for object generation in VR. In particular, we are confident that our technology-agnostic approach, which does not rely on current AI capabilities or technologies, will ensure that the results remain relevant for future systems, even if these are not constrained by the same technological limitations.
However, we acknowledge that our experiment focused on specific AI representation modes, and other variants or visualizations may yield different results. We selected these specific modes of representation because of their importance for human collaboration to explore their impact on users’ perceptions of the co-creative process and to stimulate further research and alternative implementations.
@@
Also, the limited number of possible AI modifications, which remained constant during iterative modifications, could have led participants to the conclusion that their AI partner did not actually generate object parts but retrieved them from previously modeled parts. We hypothesize that this may have had an impact on the perception of AI and its results, as there was nothing about AI’s capabilities and how its ”generative” process changed between conditions.
@@
Since we performed the study as a controlled experiment, we acknowledge limitations in external validity in favor of establishing cause-effect relationships to recognize the impacts of individual representation modes.
There are several interesting directions for future work: One important area is exploring how different methods of emphasizing the AI’s contribution process might positively impact a user’s experience with a co-creative system, including whether such methods can enhance user navigation and coordination during collaboration. Additionally, investigating various ways to incrementally visualize the AI’s generation of a 3D object could reveal how these visualizations affect users’ perceptions of the AI and its creative process. There is also a need to examine the specific characteristics of avatars for co-creative AI that may cause discomfort to users, as well as how the perceived closeness to AI agents is affected by their embodiment and which other factors might influence this perception. Finally, future research should also explore whether manipulating the visual appearance of an AI’s embodiment can alter people’s understanding and evaluation of the AI and its creations.
@@ 7. Conclusion
Research on co-creative systems has predominantly focused on how humans participate in the co-creative process, often overlooking how AI contributions are represented and perceived by users. This paper addresses this gap by examining how an AI’s contributions can be effectively represented for co-creative object generation in VR. Through a Wizard-of-Oz study in a VR-based co-creative 3D object-generating environment, we investigated the effects on the co-creative collaboration experience for three key representation modes AI can use to communicate to the user: highlighting, incremental visualization, and embodiment of AI contributions. Finally, we derive implications for the design of tools for co-creative object-building: First, designers should avoid combining highlighting with embodiment to increase user satisfaction. Second, while an embodiment increases perceived support, it also reduces ownership of the created objects. Third, designers should favor immediate visualization when the creation process is secondary.